1D-Arrays

TE[[nD Bounds (X)]] = P select (X, Bounds)

TE[[I array (X)]] = \{ \begin{matrix}l & = & P select (X, Lower); \\ u & = & P select (X, Upper); \\ s & = & − (u, l); \\ size & = & + (s, 3); \\ t & = & Allocate (size); \\ P store (t, Type, “Array”); \\ P store (t, Bounds, X); \\ In t\} \end{matrix}
TE[[P 1D select (X_{1}, X_{2})]] = \{ \begin{matrix}t_{b} & = & P select (X_{1}, Bounds); \\ l & = & P select (t_{b}, Lower); \\ t_{1} & = & − (X_{2}, l); \\ t_{2} & = & + (t_{1}, Headersize); \\ t & = & P select (X_{1}, t_{2}) \\ & In t\} \end{matrix}
TE[[P 1D store (X_{1}, X_{2}, X_{3})]] = \{ \begin{matrix}t_{b} & = & P select (X_{1}, Bounds); \\ l & = & P select (t_{b}, Lower); \\ t_{1} & = & − (X_{2}, l); \\ t_{2} & = & + (t_{1}, Headersize); \\ t & = & P store (X_{1}, t_{2}, X_{3}) \\ & In t\} \end{matrix}